# Machine Learning Interatomic Potentials Enable Molecular Dynamics Simulations of Doped MoS2

### v1

@@ Abstract
I ABSTRACT
We present the first computational framework for molecular dynamics simulation of MoS2 doped with 25 elements spanning metals, non-metals, and transition metals using Meta’s Universal Model for Atoms machine learning interatomic potential (MLIP). Benchmarking against density functional theory calculations demonstrates the accuracy of the MLIP for simulating doped-MoS2 systems and highlights opportunities for improvement. Using the MLIP, we perform heating-cooling simulations of doped-MoS2 supercells. The simulations capture complex phenomena including dopant clustering, MoS2 layer fracturing, interlayer diffusion, and chemical compound formation at orders-of-magnitude reduced computational cost compared to density functional theory. This work provides an open-source computational workflow for application-oriented design of doped-MoS2, enabling high-throughput screening of dopant candidates and optimization of compositions for targeted tribological, electronic, and optoelectronic performance. The MLIP bridges the accuracy-efficiency gap between first-principles methods and empirical potentials, and the framework offers unprecedented opportunities for large-scale materials discovery in two-dimensional doped material systems.
@@ I ABSTRACT
Two-dimensional materials have fundamentally transformed materials science and engineering over the past two decades, giving rise to a new era of atomic-scale engineering [1, 2]. Among the expansive family of 2D materials, molybdenum disulfide (MoS2) stands out due to its unique combination of mechanical, electronic, and optical properties that position it at the forefront of next-generation technological applications [3, 4, 5]. Unlike its semi-metallic counterpart graphene, which lacks an intrinsic bandgap, MoS2 exhibits a tunable bandgap that transitions from indirect (1.2 eV) in bulk form to direct (1.9 eV) in monolayer configuration due to quantum confinement effects [6, 7, 8]. MoS2 also has exceptional mechanical strength comparable to steel while possessing good chemical stability and processability [9, 10]. Further, MoS2 is a layered material. The layers of MoS2 are held together by weak van der Waals forces which allows for easy exfoliation to monolayer and facilitates low resistance to sliding between layers [11].
@@
These unique properties have made MoS2 a prime candidate for various applications. For example, The mechanical strength and low-resistance to sliding of MoS2 result in low friction and wear, essential for tribological applications [12, 11]. In electronics, field-effect transistors (FETs) based on single-layer MoS₂ exhibit very high current on/off ratios (exceeding 108) and excellent switching characteristics. [8]. MoS2 is also used for energy storage [13, 5] and as catalysts in hydrogen evolution reactions and oxygen evolution reactions [9, 14]. Optoelectronic applications leverage the direct bandgap of MoS2 monolayer to enable effective absorption and emission across the visible spectrum [6, 15, 7, 16]. Photodetectors based on monolayer and few-layer MoS2 exhibit ultra fast photo response with carrier extraction occurring on the femtosecond to millisecond timescale, high photoresponsivity and external quantum efficiencies reaching up to 7 % [16]. Light-emitting diodes and related heterostructures show strong photoluminesence at 665 nm with quantum yields approximate quantum yields approximately three orders of magnitude higher than bulk MoS₂ and direct bandgap emission suitable for display applications [7].
@@
Impurity atoms can be intentionally introduced into the MoS2 crystal lattice, i.e., dopants, to enhance its physical, chemical or electronic properties for the discussed applications. When Mo is substituted with Re, Ta, V and Tc, or Li, Na and K are absorbed on the surface, it creates an n-type semiconductor, increasing electron concentration and electrical conductivity [17]. Alternatively, doping with Ag, Au, Cu, C, P Nb, N, As and Sb in different sites create a p-type semiconductor, allowing it to conduct electricity through positive charge carriers [17, 18, 19]. Gas dopants such as molecular O, NO2 and NH3 can reversibly modify the electronic properties of MoS2, providing opportunities for tunable sensors and adaptive electronic devices [20]. Sb doping has been explored for thermoelectric applications, where the heavy atom mass contributes to reduced thermal conductivity while maintaining electrical conductivity [21]. Transition metal dopants including Co, Ni, Ru, and Fe, as well as non-metals like N, enhance the catalytic activity for hydrogen evolution reaction [22, 23, 24]. Finally, a wide variety of dopants have shown promise in tribological coatings, where dopant incorporation improves wear resistance and reduces friction [12, 11].
@@
The effect of dopants on MoS₂ properties and performance in target applications has been studied extensively using atomistic simulations. Most atomistic studies performed on doped MoS2 used density functional theory (DFT). DFT provides quantum-mechanically accurate descriptions of electronic structure, energetics, and chemical bonding, and it has been extensively used to study various chemical and electrical properties of MoS2 [25, 17, 7, 26, 11, 18, 27, 13]. However, the accuracy of DFT calculations comes as the cost of computational efficiency such that they are limited to tens to hundreds of atoms and timescales on the order of femtoseconds. This is an issue for modeling doped MoS2 since extended simulation sizes and timescales are needed to capture collective phenomenon such as dopant diffusion, clustering, interface formation, phase transition and long range strain fields - processes that fundamentally govern the functional performance of the material [28, 29, 30].
@@
To address the computational limitations of DFT, researchers have turned to molecular dynamics (MD) simulations using empirical approximations, or potentials. These approximations enable simulations of thousands to millions of atoms and time scales orders of magnitudes longer than DFT. For doped MoS2 specifically, one study performed MD simulations of Cr-doped MoS2 using newly developed CHARMM and CVFF potential parameters and showed that Cr doping significantly affects structural stability and increases hydrophobicity [31]. Other research teams developed ReaxFF parameters for Ni-doped MoS2 and Ti-doped MoS2, enabling reactive simulations of phase transitions from amorphous to crystalline structures during annealing [29, 32]. The Ni-doped potential was later used to study the effects of dopant composition on MoS2 crystallization [28]. More recently, a ReaxFF potential was developed to study C doped MoS2 systems [33]. However, empirical potentials suffer from fundamental limitations in transferability and accuracy, particularly when applied to systems or conditions significantly different from those used in their parameterization. The inherent trade-off between computational efficiency and chemical accuracy in empirical potentials becomes particularly problematic for doped systems where electronic effects determine dopant stability and property modifications.
@@
Recently, researchers have started using machine learning interatomic potentials (MLIP) to study MoS2 systems [34, 35, 36]. MLIPs are rapidly emerging as transformative approaches that bridge the accuracy-efficiency gap between DFT and MD. Recent advances include deep neural network potentials trained on DFT data that achieve near-quantum accuracy while enabling simulations of thousands of atoms over extended timescales [37]. One of the most recent and advances in the field of MLIPs is Meta’s Universal Model for Atoms (UMA), a family of machine-learning interatomic potential trained on half a billion unique 3D atomic structures spanning molecules, materials, and catalysts [38]. The UMA small and UMA medium variants utilize a novel mixture-of-linear-experts architecture, activating only a fraction of parameters for each atomic structure, which greatly enhances computational efficiency. UMA medium has 1.4 billion total parameters but activates only about 50 million per simulation, delivering higher accuracy with moderate additional cost, while UMA small has 150 million parameters with 6 million active parameters, which prioritizes speed and resource efficiency, achieving competitive accuracy on nearly all core tasks [38]. Both models are open-source and perform comparably or better than task-specific potentials, enabling large-scale, high-fidelity MD simulations across diverse chemical domains without finetuning. However, ML potentials have not yet been applied to computational studies of doped MoS2 systems.
@@
This study presents the first comprehensive molecular dynamics investigation of doped MoS2 systems spanning 25 dopant elements across the periodic table, enabled by Meta’s Universal Model for Atoms (UMA). We benchmark UMA small and UMA medium against reference DFT calculations, quantifying their accuracy and reliability for predicting energetics and relaxed structures of diverse doped MoS2 configurations. We then employ heating and cooling molecular dynamics to calculate dopant-dependent densities and atom mobility to reveal how different dopant species behave within the MoS2 lattice. Finally, based on dopant mobility and visual analysis of simulation trajectories, we classify the dopants into broader groups and provide an in-depth analysis of how different dopants behave in the MoS2 nanostructure. This work establishes a transferable computational workflow for validation of MLIPs for doped systems. It provides researchers with practical tools for screening dopant candidates, optimizing compositions, and predicting performance metrics relevant to tribological, electronic, and optoelectronic applications.
II Methods
II.1 Software and models
@@
We use Python and the Atomic Simulation Environment (ASE) [39] for optimization and job control. For machine-learning molecular dynamics (MLMD), FAIRChemCalculator (2.3.0) [38, 40] with UMA models- UMA small(uma-sm-1p1) and UMA medium (uma-m-1p1), and task OMAT are used. For the DFT calculations, Quantum ESPRESSO (QE) with the PBE functional [41] and PSLibrary 1.0.0 pseudopotentials [42], are used following standard QE references [43, 44]. No spin–orbit coupling (SOC) is enabled. Ovito is used for visualization [45].
@@
In both MLMD and DFT, the atomic positions are optimized with the Broyden Fletcher Goldfarb Shann (BFGS) algorithm with fixed cells. A force convergence threshold of $5\times 10^{-3}$ eV/Å is used. For MLMD, ASE is used to relax each input geometry with BFGS to the common force threshold. We record the final total energy $E_{\mathrm{tot}}$ for formation-energy analysis. All MLMD calculations are performed on either Nvidia A100s or L40 GPUs. The ASE calculators are assigned per-structure in a single-stage relax-and-evaluate workflow [40].
@@
The DFT calculations use a fixed cell relaxation method with spin polarization and no SOC. We use the PBE functional from PSLibrary 1.0.0 PAW/USPP files; element-to-pseudopotential mappings followed the run table in the SI [42]. We test gamma point calculation, $4\times 4\times 4$ and $6\times 6\times 4$ Monkhorst–Pack mesh to test energy convergence, Fermi–Dirac smearing, and plane-wave cutoffs of $124/843$ Ry for wavefunctions/charge density, respectively. The plane wave cuttoffs are selected from the maximum lowest suggested cutoffs multiplied by $1.2$ in the pseudopotentials throughout all dopants. Electronic thresholds are $10^{-4}$ with rmm-diis diagonalization. All relaxations use BFGS with the same $5\times 10^{-3}$ eV/Å force target as MLMD.
II.2 Dopant set and formation-energy formalism
[…]
@@ II.2 Dopant set and formation-energy formalism
We compute the neutral formation energies $E_{\mathrm{form}}$ with the Zhang–Northrup formalism [46]. For S-site substitution:
@@
Where $E_{tot}$ is total energy, $\mu$ is the chemical potential and $X$ indicates the dopant. Within each method (MLMD or DFT), we use consistent elemental references. We set $\mu_{\mathrm{S}}=\tfrac{1}{8}E(\mathrm{S}_{8})$ in a large box. We set $\mu_{\mathrm{Mo}}$ and $\mu_{X}$ from the lowest-energy elemental phase available in that method (e.g., bcc/fcc/hcp bulk, molecular box), selected per element in a fixed priority order. This one-method/one-reference scheme allows direct MLMD–DFT comparison of $E_{\mathrm{form}}$ values.
II.3 MLIP simulations
@@
Following equilibration, a heating cycle is performed. The system is heated from 300 K to 1000 K over 20 ps followed by equilibration at 1000 K for 20 ps in an NPT ensemble. Next the system is equilibrated at 1000 K for 100 ps in an NVT ensemble. Simulations are performed at 1000 K since that is below the the temperature at which MoS2 is reported to decompose [47, 48] while providing sufficient thermal energy for the dopant to diffuse and the local structure of MoS2 to respond to the mobility of the dopant. During these simulations, the mean square displacement (MSD) is calculated (details in SI Sect. S1) to quantify atomic mobility and dopant diffusion. Trajectory data are processed using custom Python scripts based on the Atomic Simulation Environment (ASE).
@@
All scripts used for performing these simulation are available on github. [49]
III Results and Discussion
III.1 Validation of UMA potentials
@@
The dopant-specific error magnitudes for both models are further detailed in the stacked bar plots of Figure 4. This analysis shows that the ML model achieves low error for many dopants. For UMA small, 19 of the 25 dopants tested have a cumulative absolute error of less than 1 eV and, for UMA medium, 11 of the 25 dopants tested have a cumulative absolute error of less than 1 eV. These errors are consistent with the defect formation energy errors observed for other universal MLIPs [50, 51]. These results indicate that, for the test system used for validation, UMA small has a better overall accuracy compared to UMA medium. Additionally, for our test systems, UMA small is almost twice as fast as UMA medium in steps per second. Due to the better performance of UMA small, we use this MLIP for the remainder of the work.
@@
From Figures 4 and 5, we consider a few cases where the MLIP does not provide accurate energy and local structure predictions. These errors are likely caused by the fact that the UMA training dataset contains neutral bulk systems and does not explicitly include point defects, which could induce errors in energy and localized structure [52]. First, small, highly electronegative dopants at Mo sites, i.e., O and N substituting Mo, have the largest local structure error. DFT shows that these dopants create strong, localized bonds and significant lattice contraction. The ML model does not capture this extreme distortion or the associated energy change. Second, alkali metals like Li or early transition metals like V substituting S led to large error, primarily in geometry. These dopants are much bigger than the S atom they replace, so DFT shows the local Mo–dopant bonds increasing in length. The MLIP partially failed to account for this expansion. Interestingly, the formation energy error for these cases is not very large (<0.3 eV) – meaning the model captures the thermodynamics of the system, even though it does not accurately predict the relaxed structure. Third, substituting S with Te is essentially alloying to form MoTe2-like local environments. DFT shows this process is favorable (negative formation energy; Figure S2), whereas the MLIP severely underestimates that favorability (predicting positive formation energy; Figure S3). Lastly, while dopants that are similar in character to the host elements (metals substituting Mo, or semi-metals substituting S) are predicted with good accuracy by the ML model, there are a few exceptions, for example, Fe at an S substitution site. The error that does occur for metals tends to be smaller and possibly due to effects like magnetism or charge state differences which are not accounted for in the model training data. Thus, there are still opportunities to improve the accuracy of the model through finetuning of the MLIP. This will be explored in a subsequent study.
III.2 Demonstration of the MLIP
The density of the doped systems calculated from the MLIP simulations at 300 K are shown in Figure 6. The densities of the simulated doped MoS2 systems are between 2.2 and 3.5 g/cm3. There is a wide range of densities for MoS2 reported in literature. For sputter deposited coatings, a density range from 1.90 to 5.29 g/cm3 has been reported [53]. Regardless, the fact that the model densities are reasonable compared to the large experimental range is encouraging for this demonstration of the MLIP simulations.
@@ III.2 Demonstration of the MLIP
The last group of dopants is the non-metals. This group consists of C, Cl, F, N, O, and Si. These dopants are more reactive than the metal dopants and form various chemical compounds with MoS2. The specific molecular species varies with dopant chemistry. In oxygen-doped systems, we observe oxidation of both Mo and S atoms, to form MoO3 and gaseous SO₂ molecules within the simulation box. Carbon dopants form extended chain structures that create interlayer linkages between MoS₂ layers along with gaseous CS2. Chlorine and fluorine dopants lead to the formation of molybdenum and sulfur halides, indicating strong halogenation reactions with the host lattice. As a representative case from this group, we analyze the N-doped MoS2 simulation, which forms Mo-S-N complexes.

### v2

@@ Abstract
Dopants are used to tune the performance of MoS2 in various applications. The effect of the dopants can be investigated using molecular dynamics simulations with machine learning interatomic potentials (MLIPs), but the accuracy of these potentials must first be evaluated. Here, we evaluate the accuracy of a recently developed MLIP, META’s Universal Model for Atoms (UMA), for 25 different MoS2 dopants spanning metals, non-metals, and transition metals in Mo substitution, S substitution, and intercalated positions by benchmarking the MLIP-predicted formation energy and the dopant-induced structural change against density functional theory calculations. Among the 25 dopant species considered, UMA predicts the formation energies and local structures of Ag, Al, Au, Cu, Ir, Na, Nb, Pd, Pt, Re, Rh, Ru, Si, Ta, Ti, and Zn doped MoS2 with higher accuracy than the other dopants tested. The MLIP is used to perform heating-cooling simulations of MoS2 supercells with all 25 dopants. These simulations capture complex phenomena including dopant clustering, MoS2 layer fracturing, interlayer diffusion, and chemical compound formation at orders-of-magnitude reduced computational cost compared to density functional theory. This work provides an open-source computational workflow for application-oriented design of doped-MoS2, enabling high-throughput screening of dopant candidates and optimization of compositions for targeted tribological, electronic, and optoelectronic performance.
Keywords
@@ Introduction
Transition metal dichalcogenides (TMDs) have fundamentally transformed materials science and engineering over the past two decades, giving rise to a new era of atomic-scale engineering [Kumar2022-zl, Ozugur-Uysal2022-eq]. Among the expansive family of TMDs, molybdenum disulfide (MoS2) stands out due to its unique combination of mechanical, electronic, and optical properties that position it at the forefront of next-generation technological applications [Alharthi2024-zb, Verma2024-ww, Zeimpekis2024-of]. Unlike its semi-metallic counterpart graphene, which lacks an intrinsic bandgap, MoS2 exhibits a tunable bandgap that transitions from indirect (1.2 eV) in bulk form to direct (1.9 eV) in monolayer configuration due to quantum confinement effects [Lu2014-mq, Gerbi2025-dh, Lembke2015-ui]. MoS2 also has mechanical strength comparable to steel while possessing good chemical stability and processability [Hemanth2021-bp, Pang2020-rz]. Further, MoS2 is a layered material. The layers of MoS2 are held together by weak van der Waals forces which allows for easy exfoliation to monolayer and facilitates low resistance to sliding between layers [Vazirisereshk2019-rc].
@@
These unique properties have made MoS2 a prime candidate for various applications. For example, The mechanical strength and low-resistance to sliding of MoS2 result in low friction and wear, essential for tribological applications [Faiyad2025-le, Vazirisereshk2019-rc]. In electronics, field-effect transistors (FETs) based on single-layer MoS2 exhibit very high current on/off ratios (exceeding 108) and good switching characteristics. [Lembke2015-ui]. MoS2 is also used for energy storage [Raza2025-dc, Zeimpekis2024-of] and as catalysts in hydrogen evolution reactions and oxygen evolution reactions [Hemanth2021-bp, Mphuthi2022-pf]. Optoelectronic applications leverage the direct bandgap of MoS2 monolayer to enable effective absorption and emission across the visible spectrum [Lu2014-mq, Timpel2021-ij, Gerbi2025-dh, Wang2018-or]. Photodetectors based on monolayer and few-layer MoS2 exhibit ultra fast photo response with carrier extraction occurring on the femtosecond to millisecond timescale, high photoresponsivity and external quantum efficiencies reaching up to 7 % [Wang2018-or]. Light-emitting diodes and related heterostructures show strong photoluminesence at 665 nm with quantum yields approximately three orders of magnitude higher than bulk MoS2 and direct bandgap emission suitable for display applications [Gerbi2025-dh].
@@
Impurity atoms can be intentionally introduced into the MoS2 crystal lattice, i.e., dopants, to enhance its physical, chemical or electronic properties for the discussed applications. When Mo is substituted with Re, Ta, V and Tc, or Li, Na and K are absorbed on the surface, it creates an n-type semiconductor, increasing electron concentration and electrical conductivity [Bae2024-cl]. Alternatively, doping with Ag, Au, Cu, C, P Nb, N, As and Sb in different sites create a p-type semiconductor, allowing it to conduct electricity through positive charge carriers [Bae2024-cl, Li2022-it, Kamruzzaman2021-eu]. Gas dopants such as molecular O, NO2 and NH3 can reversibly modify the electronic properties of MoS2, providing opportunities for tunable sensors and adaptive electronic devices [Koci2023-hu]. Sb doping has been explored for thermoelectric applications, where the heavy atom mass contributes to reduced thermal conductivity while maintaining electrical conductivity [Wu2024-tu]. Transition metal dopants including Co, Ni, Ru, and Fe, as well as non-metals like N, enhance the catalytic activity for hydrogen evolution reaction [Jiang2021-br, Nolan2023-ui, Wang2021-pw]. Finally, a wide variety of dopants have shown promise in tribological coatings, where dopant incorporation improves wear resistance and reduces friction [Faiyad2025-le, Vazirisereshk2019-rc].
@@
The effect of dopants on MoS2 properties and performance in target applications has been studied extensively using atomistic simulations. Most atomistic studies performed on doped MoS2 used density functional theory (DFT). DFT provides quantum-mechanically accurate descriptions of electronic structure, energetics, and chemical bonding, and it has been extensively used to study various chemical and electrical properties of MoS2 [Dolui2013-zl, Bae2024-cl, Gerbi2025-dh, Wang2023-tt, Vazirisereshk2019-rc, Li2022-it, Ramsey2025-vz, Raza2025-dc]. However, the accuracy of DFT calculations comes as the cost of computational efficiency such that they are limited to tens to hundreds of atoms and timescales on the order of femtoseconds. This is an issue for modeling doped MoS2 since extended simulation sizes and timescales are needed to capture collective phenomenon such as dopant diffusion, clustering, interface formation, phase transition and long range strain fields - processes that fundamentally govern the functional performance of the material [Romero-Garcia2024-zq, Mohammadtabar2023-eu, Rai2024-fj].
@@
To address the computational limitations of DFT, researchers have turned to molecular dynamics (MD) simulations using empirical approximations, or potentials. These approximations enable simulations of thousands to millions of atoms and time scales orders of magnitudes longer than DFT. For doped MoS2 specifically, one study performed MD simulations of Cr-doped MoS2 using newly developed CHARMM and CVFF potential parameters and showed that Cr doping significantly affects structural stability and increases hydrophobicity [Xing2022-sm]. Other research teams developed ReaxFF parameters for Ni-doped MoS2 and Ti-doped MoS2, enabling reactive simulations of phase transitions from amorphous to crystalline structures during annealing [Mohammadtabar2023-eu, Mao2022-ah]. The Ni-doped potential was later used to study the effects of dopant composition on MoS2 crystallization [Romero-Garcia2024-zq]. More recently, a ReaxFF potential was developed to study C doped MoS2 systems [Ponomarev2024-fw]. However, empirical potentials suffer from fundamental limitations in transferability and accuracy, particularly when applied to systems or conditions significantly different from those used in their parameterization. The inherent trade-off between computational efficiency and chemical accuracy in empirical potentials becomes particularly problematic for doped systems where electronic effects determine dopant stability and property modifications.
To address this issue, machine learning interatomic potentials (MLIP) are rapidly emerging as transformative approaches that bridge the accuracy-efficiency gap between DFT and MD. Researchers have started using MLIPs to study TMDs [Zhao2024-hn, Kety2024-zx, Flototto2025-fn, Siddiqui2025-tv, Siddiqui2024-ar, Liu2024-lw, Gu2019-zz, Marmolejo-Tejada2022-zs, Wang2019-dm]. Most of these studies focused on developing MLIPs for specific material systems, particularly, bilayers, heterostructures, and alloys of MoS2, MoSe2, WS2, and WSe2. MLIPs were used to simulate the growth process of MoS2/WS2 van der Waals heterostructures[Zhao2024-hn], the formation of amorphous MoS2 through melting-quenching[Kety2024-zx], sulfur vacancy dynamics in MoS2 monolayer[Flototto2025-fn], and twisted bilayers of MX2 (where M=Mo/W and X=S/Se)[Siddiqui2025-tv]. MLIP-based simulations have reproduced the vibrational spectrum of Mo1-xWxS2-2ySe2y quaternary alloy TMDs[Siddiqui2024-ar], determined the lattice thermal conductivity of monolayer MoS2(1-x)Se2x[Gu2019-zz] and monolayer MoS2(1-x)Se2x[Marmolejo-Tejada2022-zs], and calculated the mechanical properties of WS2 monolayer[Wang2019-dm] and MoS2/WS2 alloys[Liu2024-lw]. However, to the authors knowledge, MLIPs have not yet been developed or used for modeling doped MoS2 systems.
@@
For materials more generally, recent advances in MLIPs include deep neural network potentials trained on DFT data that achieve near-quantum accuracy[Frohlking2024-yv]. Recently, universal MLIPs trained on large DFT datasets containing elements and compounds spanning the periodic table have been developed. One such universal MLIP is META’s Universal Model for Atoms (UMA), a family of machine-learning interatomic potential trained on half a billion unique 3D atomic structures spanning molecules, materials, and catalysts [Wood2025-it]. The UMA small and UMA medium variants. UMA medium has 1.4 billion total model parameters and has a slower inference speed compared to UMA small which has 150 million total model parameters. Both models utilize a novel mixture-of-linear-experts architecture, activating only a fraction of parameters for each atomic structure, which greatly enhances computational efficiency [Wood2025-it]. Both models are open-source and perform comparably or better than task-specific potentials, enabling large-scale, high-fidelity MD simulations across diverse chemical domains without finetuning. Here, we evaluate META’s UMA MLIP for doped MoS2.
@@
This study presents a comprehensive validation of formation energy and local structure of 25 dopants in MoS2 systems using universal MLIPs. We benchmark UMA small and UMA medium against reference DFT calculations, quantifying their accuracy and reliability for predicting energetics and relaxed structures of three doped MoS2 configurations. We then run heating and cooling molecular dynamics to calculate dopant-dependent densities and atom mobility to reveal how different dopant species affect MoS2. Based on dopant mobility and visual analysis of simulation trajectories, we classify the dopants into groups and analyze how dopants in each group behave in the MoS2 nanostructure. This work establishes a transferable computational workflow for validation of MLIPs for doped TMD systems.
Methods
Software and models
@@ Methods
We use Python and the Atomic Simulation Environment (ASE) [Hjorth-Larsen2017-zj] for optimization and job control. For machine-learning molecular dynamics (MLMD), FAIRChemCalculator (2.3.0)  [Wood2025-it, UnknownUnknown-zq] with UMA models- UMA small(uma-sm-1p1) and UMA medium (uma-m-1p1), and task OMAT are used. For the DFT calculations, Quantum ESPRESSO (QE) with the PBE functional [Perdew1996-ij] and PSLibrary 1.0.0 pseudopotentials [Dal-Corso2014-lu], are used following standard QE references [Giannozzi2009-sf, Giannozzi2017-du]. No spin–orbit coupling (SOC) is enabled. Ovito is used for visualization [Stukowski2009-cu].
@@
In both MLMD and DFT, the atomic positions are optimized with the Broyden Fletcher Goldfarb Shann (BFGS) algorithm with fixed cells. A force convergence threshold of $5\times 10^{-3}$ eV/Å is used. For MLMD, ASE is used to relax each input geometry with BFGS to the common force threshold. We record the final total energy $E_{\mathrm{tot}}$ for formation-energy analysis. All MLMD calculations are performed on either Nvidia A100s or L40 GPUs. The ASE calculators are assigned per-structure in a single-stage relax-and-evaluate workflow [UnknownUnknown-zq].
@@
The DFT calculations use a fixed cell relaxation method with spin polarization and no SOC. We use the PBE functional from PSLibrary 1.0.0 PAW/USPP files. We test gamma point calculation, $4\times 4\times 4$ and $6\times 6\times 4$ Monkhorst–Pack mesh to test energy convergence, Fermi–Dirac smearing, and plane-wave cutoffs of $124/843$ Ry for wavefunctions/charge density, respectively. The plane wave cuttoffs are selected from the maximum lowest suggested cutoffs multiplied by $1.2$ in the pseudopotentials throughout all dopants. Electronic thresholds are $10^{-4}$ with rmm-diis diagonalization. All relaxations use BFGS with the same $5\times 10^{-3}$ eV/Å force target as MLMD.
Dopant set and formation-energy formalism
The dopant set is Ag, Al, Au, C, Cl, Cu, F, Fe, Ir, Li, N, Na, Nb, O, Pd, Pt, Re, Rh, Ru, Si, Ta, Te, Ti, V, and Zn. For each dopant $X$, we built three 48-atom MoS2 test structure: S-site substitution, Mo-site substitution, and intercalation between layers, as shown in Figure 1. Each structure was relaxed separately with MLMD and with DFT.
@@
We compute the neutral formation energies $E_{\mathrm{form}}$ with the Zhang–Northrup formalism [Zhang1991-on]. For S-site substitution:
@@
Where $E_{tot}$ is total energy, $\mu$ is the chemical potential and $X$ indicates the dopant. Within each method (MLMD or DFT), we use consistent elemental references. We set $\mu_{\mathrm{S}}=\tfrac{1}{8}E(\mathrm{S}_{8})$ in a large box. We set $\mu_{\mathrm{Mo}}$ and $\mu_{X}$ from the lowest-energy elemental phase available in that method (e.g., bcc/fcc/hcp bulk, molecular box), selected per element based their lowest energy phase identified with Material Projects [Jain2013-gm]. This one-method/one-reference scheme allows direct MLMD–DFT comparison of $E_{\mathrm{form}}$ values.
MLIP simulations
@@
Following equilibration, a heating cycle is performed. The system is heated from 300 K to 1000 K over 20 ps followed by equilibration at 1000 K for 20 ps in an NPT ensemble. Next the system is equilibrated at 1000 K for 100 ps in an NVT ensemble. Simulations are performed at 1000 K since that is below the the temperature at which MoS2 is reported to decompose [Brainard1969-ch, Chen2019-xu] while providing sufficient thermal energy for the dopant to diffuse and the local structure of MoS2 to respond to the mobility of the dopant. During these simulations, the mean square displacement (MSD) is calculated (details in SI Sect. S1) to quantify atomic mobility and dopant diffusion. Trajectory data are processed using custom Python scripts based on the Atomic Simulation Environment (ASE).
@@
All scripts used for performing these simulation are available on github. [FaiyadUnknown-dv]
Results and Discussion
Validation of UMA potentials
@@
The dopant-specific error magnitudes for both models are further detailed in the stacked bar plots of Figure 4. This analysis shows that the ML model achieves low error for many dopants. For UMA small, 19 of the 25 dopants tested have a cumulative absolute error of less than 1 eV and, for UMA medium, 11 of the 25 dopants tested have a cumulative absolute error of less than 1 eV. These errors are consistent with the defect formation energy errors observed for other universal MLIPs [Berger2025-ua, Broberg2023-rm]. These results indicate that, for the test system used for validation, UMA small has a better overall accuracy compared to UMA medium. Additionally, for our test systems, UMA small is almost twice as fast as UMA medium in steps per second. Due to the better performance of UMA small, we use this MLIP for the remainder of the work.
@@
From Figures 4 and 5, we consider a few cases where the MLIP does not provide accurate energy and local structure predictions. These errors are likely caused by the fact that the UMA training dataset contains neutral bulk systems and does not explicitly include point defects, which could induce errors in energy and localized structure [Barroso-Luque2024-xh]. First, small, highly electronegative dopants at Mo sites, i.e., O and N substituting Mo, have the largest local structure error. DFT shows that these dopants create strong, localized bonds and significant lattice contraction. The ML model does not capture this extreme distortion or the associated energy change. Second, alkali metals like Li or early transition metals like V substituting S led to large error, primarily in geometry. These dopants are much bigger than the S atom they replace, so DFT shows the local Mo–dopant bonds increasing in length. The MLIP partially failed to account for this expansion. Interestingly, the formation energy error for these cases is not very large (<0.3  eV) – meaning the model captures the thermodynamics of the system, even though it does not accurately predict the relaxed structure. Third, substituting S with Te is essentially alloying to form MoTe2-like local environments. DFT shows this process is favorable (negative formation energy; Figure S2), whereas the MLIP severely underestimates that favorability (predicting positive formation energy; Figure S3). Lastly, while dopants that are similar in character to the host elements (metals substituting Mo, or semi-metals substituting S) are predicted with good accuracy by the ML model, there are a few exceptions, for example, Fe at an S substitution site. The error that does occur for metals tends to be smaller and possibly due to effects like magnetism or charge state differences which are not accounted for in the model training data. Thus, there are still opportunities to improve the accuracy of the model through finetuning of the MLIP. This will be explored in a subsequent study.
The results presented in Figures 4 and 5 demonstrate that UMA small predicts the formation energies of Ag, Al, Au, C, Cu, Ir, Li, Na, Nb, Pd, Pt, Re, Rh, Ru, Si, Ta, Ti, Zn, and Pt dopants in MoS2 with formation energy errors below 0.3 eV for each of the three dopant sites. Furthermore, the structural effects of Ag, Al, Au, Cl, Cu, F, Fe, Ir, Na, Nb, Pd, Pt, Re, Rh, Ru, Si, Ta, Te, Ti, and Zn doping at the two substitution sites are reproduced with good accuracy, exhibiting deviations of less than 6% relative to DFT. When both energetic and structural accuracies are jointly considered, the dopants Ag, Al, Au, Cu, Ir, Na, Nb, Pd, Pt, Re, Rh, Ru, Si, Ta, Ti, and Zn are identified as being most reliably captured by the UMA small model.
Demonstration of the MLIP
The density of the doped systems calculated from the MLIP simulations at 300 K are shown in Figure 6. The densities of the simulated doped MoS2 systems are between 2.2 and 3.5 g/cm3. There is a wide range of densities for MoS2 reported in literature. For sputter deposited coatings, a density range from 1.90 to 5.29 g/cm3 has been reported [Babuska2025-dk]. Regardless, the fact that the model densities are reasonable compared to the large experimental range is encouraging for this demonstration of the MLIP simulations.
@@
The last group of dopants is the non-metals. This group consists of C, Cl, F, N, O, and Si. These dopants are more reactive than the metal dopants and form various chemical compounds with MoS2. The specific molecular species varies with dopant chemistry. In oxygen-doped systems, we observe oxidation of both Mo and S atoms, to form MoO3 and gaseous SO2 molecules within the simulation box. Carbon dopants form extended chain structures that create interlayer linkages between MoS2 layers along with gaseous CS2. Chlorine and fluorine dopants lead to the formation of molybdenum and sulfur halides, indicating strong halogenation reactions with the host lattice. As a representative case from this group, we analyze the N-doped MoS2 simulation, which forms Mo-S-N complexes.